the length of a rose garden is 6 feet more than the width. the garden is surrounded by a uniform pathway which is three feet wide. the area of the pathway alone is 216 square feet. what is the perimeter of the rose garden?
Answer by mananth(16949)

let the width of the rose garden be x feet
The length will be x+6 feet
..
Pathway is 3 feet wide.
So width and length will increas by 6 feet
..
x+6 will be the width of whole plot
x+6+6 will be the length of the whole plot
..
Area of whole plot - area of pathway = area of rose garden.
(x+6)(x+12)-216 = x(x+6)
x^2+18x+144-216=x^2+6x
12x=72
x= 6 the width
Length will be x+6 = 12
..
The perimeter of rose garden will be 2*6+2*12
=36 feet